# High-dimensional quantum XYZ product codes for biased noise

### v2

@@ III.3 Logical operators and code distance of 4D XYZ product code
${\pi\left(\textbf{c}_{1}^{T},\ \textbf{c}_{2}^{T},\textbf{d}_{1}^{T},\textbf{d}_{2}^{T}\right)}^{T}$
$\in\left(\mathcal{C}_{m_{1}+m_{2}}\backslash Im\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)$
Proof.
Our goal is to find out all independent logical operators, and our method is divided into three steps. First, finding out some operators that commutes with all stabilizer generators in Eq. (16). Second, excluding those operators which are stabilizers from the operators that we find in the first step, and the remaining operators are logical operators. Finally, we prove that the number of independent logical operator pairs that we find in the second step is equal to the dimension of the corresponding 4D XYZ product code that we prove in Theorem 1.
If the form of $P_{1}$ is the same with that of $X_{G_{1}^{\prime}}$, we must have
[…]
$=\textbf{\emph{x}}\otimes H_{x_{2}}^{T}\textbf{\emph{i}}+\textbf{\emph{y}}\otimes H_{z_{2}}^{T}\textbf{\emph{j}}$
$\in\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\otimes Im\left(H_{x_{2}}^{T},H_{z_{2}}^{T}\right)$
$\begin{bmatrix}X_{G_{2}}\\ X_{G_{3}}\\ Z_{G_{2}}\\ Z_{G_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{\emph{a}}_{1}^{\prime})\ &I\ &I\ &Y(\textbf{\emph{b}}_{1}^{\prime})\ &I\\ I\ &Y(\textbf{\emph{a}}_{2}^{\prime})\ &I\ &I\ &X(\textbf{\emph{b}}_{2}^{\prime})\\ Y(\textbf{\emph{c}}_{1}^{\prime})\ &Z(\textbf{\emph{d}}_{1}^{\prime})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{\emph{c}}_{2}^{\prime})\ &Y(\textbf{\emph{d}}_{2}^{\prime})\end{bmatrix}$ (36)
Let ${\pi\left({\textbf{\emph{a}}_{1}^{\prime\prime}}^{T},\ {\textbf{\emph{a}}_{2}^{\prime\prime}}^{T},{\textbf{\emph{b}}_{1}^{\prime\prime}}^{T},{\textbf{\emph{b}}_{2}^{\prime\prime}}^{T}\right)}^{T}\in\ker{\left(\left[H_{z1}^{T},H_{x1}^{T}\right]\right)}\otimes Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ and $\begin{bmatrix}X_{G_{2}}^{\prime\prime}\\ X_{G_{3}}^{\prime\prime}\end{bmatrix}=\begin{bmatrix}X(\textbf{\emph{a}}_{1}^{\prime\prime})\ &I\ &I\ &Y(\textbf{\emph{b}}_{1}^{\prime\prime})\ &I\\ I\ &Y(\textbf{\emph{a}}_{2}^{\prime\prime})\ &I\ &I\ &X(\textbf{\emph{b}}_{2}^{\prime\prime})\end{bmatrix}$. Next, we prove that $X_{G_{2}^{\prime\prime}}X_{G_{3}^{\prime\prime}}=\left[X\left({\textbf{\emph{a}}_{1}^{\prime\prime}}^{T}\right)\ Y\left({\textbf{\emph{a}}_{2}^{\prime\prime}}^{T}\right)\ I\ Y\left({\textbf{\emph{b}}_{1}^{\prime\prime}}^{T}\right)\ X\left({\textbf{\emph{b}}_{2}^{\prime\prime}}^{T}\right)\right]$ is a stabilizer, which can only be generated by the product of some stabilizer generators in $S$ and $V$, thus should be excluded. Considering an operator $P_{2}$ which is the product of some stabilizer generators in $S$ and $V$, it must have the following form
$P_{2}$ $=\left[\textbf{\emph{s}}^{T},\textbf{\emph{t}}^{T}\right]\begin{bmatrix}X^{\left(I_{m_{1}}\otimes H_{x_{2}}^{T}\right)}&Y^{\left(I_{m_{1}}\otimes H_{z_{2}}^{T}\right)}&Z^{\left(H_{z_{1}}\otimes I_{n_{B}}\right)}&I&I\\ I&I&Z^{\left(H_{x1}\otimes I_{n_{B}}\right)}&Y^{\left(I_{m_{2}}\otimes H_{x_{2}}^{T}\right)}&X^{\left(I_{m_{2}}\otimes H_{z_{2}}^{T}\right)}\end{bmatrix}$ (37)
If the form of $P_{2}$ is the same with $X_{G_{2}^{\prime\prime}}X_{G_{3}^{\prime\prime}}$, we must have
Thus, we have $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}\in\ker\left(\left[H_{z1}^{T}\otimes I_{n_{B}},H_{x1}^{T}\otimes I_{n_{B}}\right]\right)=\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\mathcal{C}_{n_{B}}$. Let $\begin{bmatrix}\textbf{\emph{s}}\\ \textbf{\emph{v}}\end{bmatrix}=\begin{bmatrix}\textbf{\emph{x}}\\ \textbf{\emph{y}}\end{bmatrix}\otimes\textbf{\emph{i}}$, where $\begin{bmatrix}\textbf{\emph{x}}\\ \textbf{\emph{y}}\end{bmatrix}=\left[x_{1},\cdots,x_{m_{1}},y_{1},\cdots,y_{m_{y}}\right]^{T}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}$ and $\textbf{\emph{i}}\in\mathcal{C}_{n_{B}}$, we have
$\begin{bmatrix}I_{m_{1}}\otimes H_{x_{2}}&\textbf{0}\\ I_{m_{1}}\otimes H_{z_{2}}&\textbf{0}\\ \textbf{0}&I_{m_{2}}\otimes H_{x_{2}}\\ \textbf{0}&I_{m_{2}}\otimes H_{z_{2}}\end{bmatrix}\begin{bmatrix}\textbf{\emph{x}}\\ \textbf{\emph{y}}\end{bmatrix}\otimes\textbf{\emph{i}}=\begin{bmatrix}\textbf{\emph{x}}\otimes H_{x_{2}}\textbf{\emph{i}}\\ \textbf{\emph{x}}\otimes H_{z_{2}}\textbf{\emph{i}}\\ \textbf{\emph{y}}\otimes H_{x_{2}}\textbf{\emph{i}}\\ \textbf{\emph{y}}\otimes H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}$ (39)
$=\begin{bmatrix}x_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{x_{2}}\textbf{\emph{i}}\\ x_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{z_{2}}\textbf{\emph{i}}\\ y_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{x_{2}}\textbf{\emph{i}}\\ y_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}$
It can be seen that
$\pi\left(\begin{bmatrix}x_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{x_{2}}\textbf{\emph{i}}\\ x_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{z_{2}}\textbf{\emph{i}}\\ y_{1}H_{x_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{x_{2}}\textbf{\emph{i}}\\ y_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}\right)=\begin{bmatrix}x_{1}H_{x_{2}}\textbf{\emph{i}}\\ x_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ x_{m_{1}}H_{x_{2}}\textbf{\emph{i}}\\ x_{m_{1}}H_{z_{2}}\textbf{\emph{i}}\\ y_{1}H_{x_{2}}\textbf{\emph{i}}\\ y_{1}H_{z_{2}}\textbf{\emph{i}}\\ \vdots\\ y_{m_{2}}H_{x_{2}}\textbf{\emph{i}}\\ y_{m_{2}}H_{z_{2}}\textbf{\emph{i}}\end{bmatrix}$ (40)
$\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\otimes Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}$
which means $X_{G_{2}^{\prime\prime}}X_{G_{3}^{\prime\prime}}$ is a stabilizer. Thus, for vectors ${\pi\left(\textbf{\emph{a}}_{1}^{T},\textbf{\emph{a}}_{2}^{T},\textbf{\emph{b}}_{1}^{T},\textbf{\emph{b}}_{2}^{T}\right)}^{T}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$, the corresponding operators $\begin{bmatrix}X_{L_{2}}\\ X_{L_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{\emph{a}}_{1})\ &I\ &I\ &Y(\textbf{\emph{b}}_{1})\ &I\\ I\ &Y(\textbf{\emph{a}}_{2})\ &I\ &I\ &X(\textbf{\emph{b}}_{2})\end{bmatrix}$ are logical operators.
Similarly, for vector $\left({\textbf{\emph{c}}_{1}^{\prime}}^{T},{\textbf{\emph{d}}_{1}^{\prime}}^{T}\right)^{T}\in\ker{\left(I_{m_{1}}\otimes H_{x_{2}}^{T},I_{m_{1}}\otimes H_{z2}^{T}\right)}$, $\left({\textbf{\emph{c}}_{2}^{\prime}}^{T},\ {\textbf{\emph{d}}_{2}^{\prime}}^{T}\right)^{T}\in\ker{\left(I_{m_{2}}\otimes H_{x2}^{T},I_{m_{2}}\otimes H_{z2}^{T}\right)}$ and $\begin{bmatrix}Z_{G_{2}}^{\prime\prime}\\ Z_{G_{3}}^{\prime\prime}\end{bmatrix}=\begin{bmatrix}Y(\textbf{\emph{c}}_{1}^{\prime\prime})\ &Z(\textbf{\emph{d}}_{1}^{\prime\prime})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{\emph{c}}_{2}^{\prime\prime})\ &Y(\textbf{\emph{d}}_{2}^{\prime\prime})\end{bmatrix}$, the operator $Z_{G_{2}^{\prime\prime}}Z_{G_{3}^{\prime\prime}}=\left[Y\left({\textbf{\emph{c}}_{1}^{\prime\prime}}^{T}\right)\ \ Z\left({\textbf{\emph{c}}_{2}^{\prime\prime}}^{T}\right)\ \ I\ \ Z\left({\textbf{\emph{d}}_{1}^{\prime\prime}}^{T}\right)\ \ Y\left({\textbf{\emph{d}}_{2}^{\prime\prime}}^{T}\right)\right]$ is a stabilizer which can only be generated by the product of some stabilizer generators in $T$ and $U$ and thus should be excluded.
To sum up, for vectors ${\pi\left(\textbf{\emph{a}}_{1}^{T},\ \textbf{\emph{a}}_{2}^{T},\textbf{\emph{b}}_{1}^{T},\textbf{\emph{b}}_{2}^{T}\right)}^{T}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ and ${\pi\left(\textbf{\emph{c}}_{1}^{T},\ \textbf{\emph{c}}_{2}^{T},\textbf{\emph{d}}_{1}^{T},\textbf{\emph{d}}_{2}^{T}\right)}^{T}\in\left(\mathcal{C}_{m_{1}+m_{2}}\backslash Im\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)$, the corresponding $\begin{bmatrix}X_{L_{2}}\\ X_{L_{3}}\end{bmatrix}=\begin{bmatrix}X(\textbf{\emph{a}}_{1})\ &I\ &I\ &Y(\textbf{\emph{b}}_{1})\ &I\\ I\ &Y(\textbf{\emph{a}}_{2})\ &I\ &I\ &X(\textbf{\emph{b}}_{2})\end{bmatrix}$ and $\begin{bmatrix}Z_{L_{2}}\\ Z_{L_{3}}\end{bmatrix}=\begin{bmatrix}Y(\textbf{\emph{c}}_{1})\ &Z(\textbf{\emph{d}}_{1})\ &I\ &I\ &I\\ I\ &I\ &I\ &Z(\textbf{\emph{c}}_{2})\ &Y(\textbf{\emph{d}}_{2})\end{bmatrix}$ must be logical operators.
Notice that the dimension of vector space $\mathcal{C}_{m_{3}+m_{4}}\backslash Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}$ is $m_{3}+m_{4}-\dim\left(Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)=m_{3}+m_{4}-dim\left(row\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)=\dim\left(\ker\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$. Similarly, the dimension of vector space $\mathcal{C}_{m_{1}+m_{1}}\backslash Im\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}$ is $m_{1}+m_{2}-\dim\left(Im\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)=m_{1}+m_{2}-dim\left(row\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)\right)=\dim\left(\ker\left(\left[H_{x_{1}}^{T},H_{z_{1}}^{T}\right]\right)\right)$. Thus, the total number of independent logical operators $X_{L_{2}}$ and $X_{L_{3}}$ and that of $Z_{L_{2}}$ and $Z_{L_{3}}$ are both $\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)\right)$. Moreover, for any vector $\textbf{\emph{r}}\in\ker\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$, we can find a vector $\textbf{\emph{w}}\in\left(\mathcal{C}_{m_{1}+m_{2}}\backslash Im\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\otimes\ker\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)$, such that $\textbf{\emph{r}}\cdot\textbf{\emph{w}}^{T}=1$, which means the corresponding $X_{L}$ and $Z_{L}$ anti-commute. Thus, there are $\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)$ pairs of the second type logical operators.
The total number of the first and the second types of logical operators is $\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)+\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)$. Now, we prove that this number is equal to the dimension of the corresponding 4D XYZ product code as we prove in Theorem 1.
$\dim\left(\ker\left(\begin{bmatrix}H_{x_{1}}\\ H_{z_{1}}\end{bmatrix}\right)\right)\times\dim\left(\ker\left(\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)\right)$ (41)
$+\dim\left(\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\right)\times\dim\left(\ker{\left(\left[H_{z_{2}}^{T},H_{x_{2}}^{T}\right]\right)}\right)$
$=a\left[b-\left(n_{B}-m_{3}-m_{4}\right)\right]+b\left[a-\left(n_{A}-m_{1}-m_{2}\right)\right]$
$+\left(n_{A}-m_{1}-m_{2}\right)\left(n_{B}-m_{3}-m_{4}\right)$
which is equal to the dimension as we prove in Theorem 1. Thus, the proof is completed. ∎
@@
Next, we will prove the upper bound of code distance of 4D XYZ product code in corollary 4. Before proving it, we give out Lemma 2 which will be used in the proof of corollary 4.
@@
Giving a vector space $\mathcal{C}_{n}$ with dimension $n$ over $GF(2)$, two binary matrices $H_{1}$ and $H_{2}$ with size $n\times m_{1}$ and $m_{2}\times n$ respectively, the minimum weight of non-zero vector in $\mathcal{C}_{n}\backslash Im(H_{1})$ and $\mathcal{C}_{n}\backslash\ker(H_{2})$ is one.
The proof of Lemma 2 is given in the Appendix A.
@@
Secondly, considering the second type of logical operators, and the upper bound of the minimum weight of vector ${\pi\left(\textbf{\emph{a}}_{1}^{T},\textbf{\emph{a}}_{2}^{T},\textbf{\emph{b}}_{1}^{T},\textbf{\emph{b}}_{2}^{T}\right)}^{T}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is $d_{3}$. Observing Eq. (30), it is obvious that the minimum weight of the logical operators $X_{L_{2}}$ and $X_{L_{3}}$ is no more than $d_{3}$. Similarly, the the minimum weight of the logical operators $Z_{L_{2}}$ and $Z_{L_{3}}$ is no more than $d_{4}$.
@@
Here we exploit Monte Carlo method proposed in Ref. [19, 20] to determine code distance of 4D Chamon codes and 4D XYZ concatenated codes and the results are consistent with Corollary 4. Further, we compare the dimension and code distance between 4D Chamon codes, 4D toric codes, 4D XYZ concatenated codes and 4D homological concatenated codes in Table 1, where $n_{1},n_{2},n_{3},n_{4}$ are the code length of the repetition codes. This comparison results support that using the same two CSS codes, the non-CSS code constructed by 4D XYZ product may have higher dimension or code distance than those of the CSS code constructed by 4D homological product.
Table 1: Comparison of 4D Chamon codes, 4D toric codes, 4D XYZ concatenated codes and 4D homological concatenated codes in dimension and code distance.
$n_{1}$, $n_{2}$, $n_{3}$, $n_{4}$ dimension code distance
@@
IV Numerical simulations
In this section, we exploit FDBP-OSD-0 proposed in Ref. [16] to study the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ concatenated codes under the independent single-qubit Pauli noise model, in which each qubit independently suffers a Pauli error $I$, $X$, $Y$, or $Z$ with error probability $1-p$, $p_{x}$, $p_{y}$, $p_{z}$, respectively, where $p=p_{x}+p_{y}+p_{z}$ is physical qubit error rate. Under this noise model, the bias rate of Pauli $Z$ noise is defined as

### v3

@@ 3.3 Logical operators and code distance of 4D XYZ product code
${\pi\left(\textbf{c}_{1}^{T},\ \textbf{d}_{1}^{T},\textbf{c}_{2}^{T},\textbf{d}_{2}^{T}\right)}^{T}$
$\in\left(\mathcal{C}_{m_{1}+m_{2}}\backslash Im\left(\begin{bmatrix}H_{z_{1}}\\ H_{x_{1}}\end{bmatrix}\right)\right)\otimes\ker\left(\left[H_{x_{2}}^{T},H_{z_{2}}^{T}\right]\right)$
The proof of Theorem 2 is given in the Appendix A.
@@
Next, we will prove the upper bound of code distance of 4D XYZ product code in corollary 4. Before proving it, we first present Lemma 2 which will be used later.
@@
Given a vector space $\mathcal{C}_{n}$ with dimension $n$ over $GF(2)$, two binary matrices $H_{1}$ and $H_{2}$ with size $n\times m_{1}$ and $m_{2}\times n$ respectively, the minimum weight of non-zero vector in $\mathcal{C}_{n}\backslash Im(H_{1})$ and $\mathcal{C}_{n}\backslash\ker(H_{2})$ is one.
The proof of Lemma 2 is given in the Appendix B.
@@
For the second type of logical operators, the upper bound of the minimum weight for the vector ${\pi\left(\textbf{\emph{a}}_{1}^{T},\textbf{\emph{a}}_{2}^{T},\textbf{\emph{b}}_{1}^{T},\textbf{\emph{b}}_{2}^{T}\right)}^{T}\in\ker{\left(\left[H_{z_{1}}^{T},H_{x_{1}}^{T}\right]\right)}\otimes\left(\mathcal{C}_{m_{3}+m_{4}}\backslash Im\begin{bmatrix}H_{x_{2}}\\ H_{z_{2}}\end{bmatrix}\right)$ is $d_{3}$. Observing Eq. (31), it is obvious that the minimum weight of the logical operators $X_{L_{2}}$ and $X_{L_{3}}$ is no more than $d_{3}$. Similarly, the the minimum weight of the logical operators $Z_{L_{2}}$ and $Z_{L_{3}}$ is no more than $d_{4}$.
@@
Here we exploit Monte Carlo method proposed in Ref. [21, 22] to verify the code distance of 4D Chamon codes and 4D XYZ product concatenated codes and the results are consistent with Corollary 4. Further, we compare the code dimension and the code distance between 4D Chamon codes, 4D toric codes, 4D XYZ product concatenated codes and 4D homological product concatenated codes in Table 1, where $n_{1},n_{2},n_{3},n_{4}$ are the code length of the repetition codes. The comparison results support that using the same two CSS codes, 4D XYZ product can construct non-CSS codes with higher code dimension or code distance than the CSS codes constructed by 4D homological product.
Table 1: Comparison of 4D Chamon codes, 4D toric codes, 4D XYZ product concatenated codes and 4D homological product concatenated codes in code dimension and code distance.
$n_{1}$, $n_{2}$, $n_{3}$, $n_{4}$ code dimension code distance
@@
4 Numerical simulations
In this section, we exploit FDBP-OSD-0 proposed in Ref. [18] to study the error-correcting performance of 3D Chamon codes, 4D Chamon codes and 4D XYZ product concatenated codes under the independent single-qubit Pauli noise model, in which each qubit independently suffers a Pauli error $I$, $X$, $Y$, or $Z$ with error probability $1-p$, $p_{x}$, $p_{y}$, $p_{z}$, respectively, where $p=p_{x}+p_{y}+p_{z}$ is physical qubit error rate. Under this noise model, the bias rate of Pauli $Z$ noise is defined as